# Invariant subspaces and PCA in nearly matrix multiplication time

### v3

@@ Appendix G Deflation and bit complexity of PCA
Since its introduction in the early twentieth century [Pea01, Hot33], Principal Component Analysis is one of the most important tools in statistics, data science, and machine learning. It can be used, for example, to visualize data, to reduce dimensionality, or to remove noise in measurements; cf. [Jol02, DK96] for reviews on the vast bibliography. In its simplest formulation, given a (centered) data matrix $\boldsymbol{\mathrm{X}}\in\mathbb{R}^{m\times n}$, the goal is to find a $k$-dimensional embedding $\boldsymbol{\mathrm{C}}_{k}$, where $k<n$, that maximizes the sample variance, which can be written as an optimization problem
@@ G.1 Vanilla PCA
With this short introduction, we can now describe how to enhance the analysis of PCA algorithms by using our results for spectral projectors to obtain forward-error guarantees for all the SVD-related computations. Using our Theorem 1.1, the $\mathsf{SIGMAK}$ algorithm of Proposition E.2, and the $\mathsf{DEFLATE}$ algorithm of [BKMS21], as described in Algorithm 7, we can state the following result for the classic (“vanilla”) PCA, which provides forward error guarantees in matrix multiplication time regardless how large $k$ is. This is can serve as the backbone for more advanced algorithms that are analyzed in the next section, which typically first use random projections to reduce the matrix size to approximately $O(k\times k)$ and then perform the SVD computation on the smaller matrix.
@@ G.2 Low-rank PCA
In many applications of PCA the parameter $k$ is chosen to be small, i.e. $k\ll d$. This has led to an extensive research area on the so-called low-rank approximation algorithms for PCA. One of the earliest works which introduced this type of randomized low-rank approximations is [FKV98, FKV04]. Some other landmark works in the field include the analysis of randomized PCA and low-rank approximation [MRT11, HMT11, Sar06, CW09, CW17], and the pioneering Block-Krylov PCA of [MM15], which is essentially optimal in the matrix-vector query model [SEAR18] . The approximation accuracy of low-rank approximation-based PCA methods is often measured with respect to the spectral or the Frobenius norm error, i.e. the matrix $\widetilde{\boldsymbol{\mathrm{C}}}_{k}$ that is returned should satisfy $\boldsymbol{\mathrm{C}}_{k}^{\top}\boldsymbol{\mathrm{C}}_{k}=\boldsymbol{\mathrm{I}}_{k\times k}$ and:
@@
Most of the aforementioned low-rank approximation type algorithms assume exact arithmetic. A subtle detail is that they often rely on the computation of an (exact) SVD of some smaller submatrix, which is not realistic due to the Abel-Ruffini theorem. This can be justified for practical reasons since SVD can be approximated in polynomial time to arbitrary accuracy using classical solvers [GVL13]. Quoting [HMT11]:
@@
Assuming an exact SVD algorithm, the seminal analysis of the Block-Krylov PCA of [MM15] can in fact provide per-singular vector guarantees, which are much stronger than the classical norm-wise bounds. If $O(T_{SVD}(n))$ is the time required for the SVD for a (square) matrix of size $n$, the arithmetic complexity of Block-Krylov PCA is
$O\left({\tt nnz}(A)\tfrac{k\log(n)}{\sqrt{\epsilon}}+T_{SVD}(\tfrac{k\log(n)}{\sqrt{\epsilon}})\right).$
The complexity is favorable for coarse accuracy $\epsilon$ (i.e. when $\epsilon=\Theta(1)$) and small rank $k$. It is not suitable, however, for larger $k$. E.g., it can be the case in applications where we need to keep $k=n/20$ of the original dimensions. The same holds when higher accuracy is required, e.g. when $\epsilon=1/\operatorname{poly}(n)$. Then the complexity of these methods is already higher than standard eigensolvers.
In finite precision the landscape is even less clear. The only work related to low-rank approximation PCA that we are aware of with end-to-end bit complexity bounds is [MMS18] for the approximation of matrix functions applied on vectors. They prove that the Lanczos method can be stably applied to compute a vector $\boldsymbol{\mathrm{u}}$ such that the quantity $\|\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{u}}\|$ approximates $\|\boldsymbol{\mathrm{A}}\|$. This is a backward-approximate solution for the top-1 singular vector.
For general invariant subspaces $(k>1)$, consider one of the simplest randomized low-rank approximation algorithms, often referred to as “subspace iteration” [HMT11] or “simultaneous iteration” [MM15]. The algorithm first samples a matrix $\boldsymbol{\mathrm{G}}\in\mathbb{R}^{n\times l}$ with i.i.d standard normal elements, where $l=\Theta(k)$, and computes $\boldsymbol{\mathrm{Y}}\leftarrow(\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{T})^{q}\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{G}}$, where typically $q\approx\log(\min\{m,n\})$. It then returns the $\boldsymbol{\mathrm{Q}}$ factor from the economy QR factorization of $\boldsymbol{\mathrm{Y}}$. It can be shown (see e.g. Corollary 10.10 of [HMT11]) that
@@
The returned matrix has slightly more than $k$ columns, often referred to as “oversampling,” but, importantly, there is no explicit SVD involved in the computation. Even in this case, floating point arithmetic already spoils the approximation guarantees: we cannot compute $\boldsymbol{\mathrm{Q}}$ without rounding errors. Stability analysis of QR factorization is typically carried out in the backward-error sense [Hig02, DDH07], which makes the analysis of low-rank approximation PCA algorithms even more complicated.
@@ G.3 Bit complexity analysis of Block-Krylov PCA
We now analyze the seminal Block-Krylov iteration algorithm of [MM15], which is listed in Algorithm 8 for convenience (in exact arithmetic). For the floating point analysis the following methodology is used. The main result is stated in Corollary 4.1.
Algorithm 8 Block-Krylov Iteration (Alg. 2 of [MM15])
@@
Given $\boldsymbol{\mathrm{G}}$ and $\boldsymbol{\mathrm{X}}$, we can construct the Krylov matrix $\boldsymbol{\mathrm{K}}$ using $O(qk)$ calls to $\mathsf{MM}$. The first call $\boldsymbol{\mathrm{X}}_{1}\leftarrow\mathsf{MM}(\boldsymbol{\mathrm{X}},\boldsymbol{\mathrm{G}})$ returns a matrix $\boldsymbol{\mathrm{X}}_{1}=\boldsymbol{\mathrm{X}}\boldsymbol{\mathrm{G}}+\boldsymbol{\mathrm{E}}_{1}$ where $\|\boldsymbol{\mathrm{E}}_{1}\|\leq O(\operatorname{poly}(mnk)\textbf{{u}})\|\boldsymbol{\mathrm{X}}\|\|\boldsymbol{\mathrm{G}}\|\leq O(\operatorname{poly}(mnk)\textbf{{u}})$, since we assumed unit norms. If we keep performing multiplications recursively to build the Krylov matrix $\|\widetilde{\boldsymbol{\mathrm{K}}}\|$, after $O(q)$ multiplications it holds that $\|\widetilde{\boldsymbol{\mathrm{K}}}-\boldsymbol{\mathrm{K}}\|\leq O(\operatorname{poly}(mnkq)\textbf{{u}})$. The total cost is $O(qmnk^{\omega-2})$ floating point operations.
@@
So far we have approximated $\boldsymbol{\mathrm{K}}$ by $\widetilde{\boldsymbol{\mathrm{K}}}=\boldsymbol{\mathrm{K}}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{K}}}$. In the following steps we will need an approximation for the condition number of $\boldsymbol{\mathrm{K}}$. We can get such an approximation using a variant of the $\mathsf{COND}$ algorithm (Corollary E.1), which internally uses the $\mathsf{SIGMAK}$ algorithm of Proposition E.2 based on counting queries to approximate the condition number. In that case, we start with some arbitrary error $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{K}}}\|$, and we keep keep halving it until we have a sufficiently good approximation. The algorithm executes $O(qmnk^{\omega-2}\operatorname{polylog}(qmnk\kappa(\boldsymbol{\mathrm{K}})))$ floating point operations using $O(\operatorname{polylog}(qmnk\kappa(\boldsymbol{\mathrm{K}})))$ bits, and returns a value $\widetilde{\kappa}\in\Theta(\kappa(\boldsymbol{\mathrm{K}}))$.
@@
$\textbf{{u}}\leq\epsilon^{\prime}\frac{c}{\operatorname{poly}(mnqk)\kappa(\boldsymbol{\mathrm{K}})},$
@@
Let $\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}=\boldsymbol{\mathrm{E}}_{1}\boldsymbol{\mathrm{\Phi}}_{2}+\boldsymbol{\mathrm{E}}_{\mathsf{QR}}$. We have that $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}\|=\|\boldsymbol{\mathrm{E}}_{1}\boldsymbol{\mathrm{\Phi}}_{2}+\boldsymbol{\mathrm{E}}_{\mathsf{QR}}\|\leq\epsilon^{\prime}+\boldsymbol{\mathrm{E}}_{\mathsf{QR}}.$ From Theorem A.6 we can ensure that $\|\boldsymbol{\mathrm{E}}_{\mathsf{QR}}\|\leq\epsilon^{\prime}$ if we use $O(\log(\tfrac{mnqk\kappa(\widetilde{\boldsymbol{\mathrm{K}}})}{\epsilon^{\prime}}))$ bits. Thus, $\widetilde{\boldsymbol{\mathrm{Q}}}$ is an approximate orthogonal basis for the range of the true Block-Krylov matrix $\boldsymbol{\mathrm{K}}$.
To summarize, as long as $\textbf{{u}}\leq\epsilon^{\prime}\frac{c}{\operatorname{poly}(mnqk)\kappa(\boldsymbol{\mathrm{K}})}$, which can be achieved by replacing $\kappa(\boldsymbol{\mathrm{K}})$ with $\widetilde{\kappa}$ above, then we can compute $\widetilde{\boldsymbol{\mathrm{Q}}}=\boldsymbol{\mathrm{Q}}\boldsymbol{\mathrm{\Phi}}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}$, where $\boldsymbol{\mathrm{Q}}$ is the true orthonormal basis from the QR factorization of $\boldsymbol{\mathrm{K}}$, $\boldsymbol{\mathrm{\Phi}}$ is a $qk\times qk$ orthogonal matrix, and $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}\|\leq\epsilon^{\prime}$. The total cost is $O(m(qk)^{\omega-1})$ floating point operations for Theorem A.6, using $O\left(\log(\tfrac{mnqk\kappa(\boldsymbol{\mathrm{K}})}{\epsilon^{\prime}})\right)$ bits.
@@
In Step 5, Algorithm 8 forms the matrix $\boldsymbol{\mathrm{M}}$ to compute its top-$k$ singular vectors. To analyze the computation of $\boldsymbol{\mathrm{M}}$, the first observation is that $\boldsymbol{\mathrm{Q}}$ in lines 4-7 can be replaced by any basis for the column space of $\boldsymbol{\mathrm{K}}$. In particular, we replace it by $\widetilde{\boldsymbol{\mathrm{Q}}}$. We then perform the multiplication in two steps: $\widetilde{\boldsymbol{\mathrm{M}}}_{1}=\mathsf{MM}(\widetilde{\boldsymbol{\mathrm{Q}}}^{\top},\boldsymbol{\mathrm{X}})$, where $\widetilde{\boldsymbol{\mathrm{M}}}_{1}=\widetilde{\boldsymbol{\mathrm{Q}}}^{\top}\boldsymbol{\mathrm{X}}+\boldsymbol{\mathrm{E}}_{1}^{\mathsf{MM}}$ where $\|\boldsymbol{\mathrm{E}}_{1}^{\mathsf{MM}}\|\leq O(\operatorname{poly}(qknm)\textbf{{u}})\|\widetilde{\boldsymbol{\mathrm{Q}}}\|\|\boldsymbol{\mathrm{X}}\|\in O(\operatorname{poly}(qknm)\textbf{{u}})$. Then we compute $\widetilde{\boldsymbol{\mathrm{M}}}\leftarrow\mathsf{MM}(\widetilde{\boldsymbol{\mathrm{M}}}_{1},\widetilde{\boldsymbol{\mathrm{M}}}_{1}^{\top})=\widetilde{\boldsymbol{\mathrm{M}}}_{1}\widetilde{\boldsymbol{\mathrm{M}}}_{1}^{\top}+\boldsymbol{\mathrm{E}}_{2}^{\mathsf{MM}}$ where $\|\boldsymbol{\mathrm{E}}_{2}^{\mathsf{MM}}\|\leq O(\operatorname{poly}(qkn)\textbf{{u}})\|\boldsymbol{\mathrm{M}}_{1}\|^{2}\in O(\operatorname{poly}(qkn)\textbf{{u}})$, where the last is implied if u is sufficiently smaller than $1/\operatorname{poly}(qmnk)$. Putting everything together we can write
@@
where $\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}$ contains the errors in $\widetilde{\boldsymbol{\mathrm{Q}}}$ and also the ones from the multiplication in floating point, and $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|\leq O(\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}\|)\in O(\epsilon^{\prime})$, where $\epsilon^{\prime}$ is as in Step 3. So far we have assumed $O(\log(\tfrac{mnqk\kappa(\widetilde{\boldsymbol{\mathrm{K}}})}{\epsilon^{\prime}}))$ bits.

### v4

@@ Appendix G Deflation and bit complexity of PCA
Since its introduction in the early twentieth century [116, 72], Principal Component Analysis is one of the most important tools in statistics, data science, and machine learning. It can be used, for example, to visualize data, to reduce dimensionality, or to remove noise in measurements; cf. [79, 45] for reviews on the vast bibliography. In its simplest formulation, given a (centered) data matrix $\boldsymbol{\mathrm{X}}\in\mathbb{R}^{m\times n}$, the goal is to find a $k$-dimensional embedding $\boldsymbol{\mathrm{C}}_{k}$, where $k<n$, that maximizes the sample variance, which can be written as an optimization problem
@@ G.1 Vanilla PCA
With this short introduction, we can now describe how to enhance the analysis of PCA algorithms by using our results for spectral projectors to obtain forward-error guarantees for all the SVD-related computations. Using our Theorem 1.1, the $\mathsf{SIGMAK}$ algorithm of Proposition E.2, and the $\mathsf{DEFLATE}$ algorithm of [19], as described in Algorithm 7, we can state the following result for the classic (“vanilla”) PCA, which provides forward error guarantees in matrix multiplication time regardless how large $k$ is. This is can serve as the backbone for more advanced algorithms that are analyzed in the next section, which typically first use random projections to reduce the matrix size to approximately $O(k\times k)$ and then perform the SVD computation on the smaller matrix.
@@ G.2 Low-rank PCA
In many applications of PCA the parameter $k$ is chosen to be small, i.e. $k\ll d$. This has led to an extensive research area on the so-called low-rank approximation algorithms for PCA. One of the earliest works which introduced this type of randomized low-rank approximations is [53, 54]. Some other landmark works in the field include the analysis of randomized PCA and low-rank approximation [100, 67, 120, 33, 34], and the pioneering Block-Krylov PCA of [106], which is essentially optimal in the matrix-vector query model [125]. The approximation accuracy of low-rank approximation-based PCA methods is often measured with respect to the spectral or the Frobenius norm error, i.e. the matrix $\widetilde{\boldsymbol{\mathrm{C}}}_{k}$ that is returned should satisfy $\boldsymbol{\mathrm{C}}_{k}^{\top}\boldsymbol{\mathrm{C}}_{k}=\boldsymbol{\mathrm{I}}_{k\times k}$ and:
@@
Most of the aforementioned low-rank approximation type algorithms assume exact arithmetic. A subtle detail is that they often rely on the computation of an (exact) SVD of some smaller submatrix, which is not realistic due to the Abel-Ruffini theorem. This can be justified for practical reasons since SVD can be approximated in polynomial time to arbitrary accuracy using classical solvers [64]. Quoting [67]:
@@
Assuming an exact SVD algorithm, the seminal analysis of the Block-Krylov PCA of [106] can in fact provide per-singular vector guarantees, which are much stronger than the classical norm-wise bounds. Then arithmetic complexity of Block-Krylov PCA depends on $\operatorname{poly}(\tfrac{k}{\sqrt{\epsilon}})$. It is favorable for coarse accuracy $\epsilon$ (i.e. when $\epsilon=\Theta(1)$) and small rank $k$. It is not suitable, however, for larger $k$. E.g., it can be the case in applications where we need to keep $k=n/20$ of the original dimensions. The same holds when higher accuracy is required, e.g. when $\epsilon=1/\operatorname{poly}(n)$. Then the complexity of these methods is already higher than standard eigensolvers.
In finite precision, the landscape is even less clear. At the time of this submission, the only work related to low-rank approximation PCA that we are aware of with end-to-end bit complexity bounds is [107] for the approximation of matrix functions applied on vectors. They prove that the Lanczos method can be stably applied to compute a vector $\boldsymbol{\mathrm{u}}$ such that the quantity $\|\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{u}}\|$ approximates $\|\boldsymbol{\mathrm{A}}\|$. This is a backward-approximate solution for the top-1 singular vector. Concurrently with this work, and independently, [80] analyzed the bit complexity of Block-Krylov PCA and achieved similar bounds as ours, albeit with different techniques. However, [80] did not describe how to compute the condition number which is required in order to adjust the machine precision. Our analysis covers the computation of all the involved parameters.
For general invariant subspaces $(k>1)$, consider one of the simplest randomized low-rank approximation algorithms, often referred to as “subspace iteration” [67] or “simultaneous iteration” [106]. The algorithm first samples a matrix $\boldsymbol{\mathrm{G}}\in\mathbb{R}^{n\times l}$ with i.i.d standard normal elements, where $l=\Theta(k)$, and computes $\boldsymbol{\mathrm{Y}}\leftarrow(\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{A}}^{T})^{q}\boldsymbol{\mathrm{A}}\boldsymbol{\mathrm{G}}$, where typically $q\approx\log(\min\{m,n\})$. It then returns the $\boldsymbol{\mathrm{Q}}$ factor from the economy QR factorization of $\boldsymbol{\mathrm{Y}}$. It can be shown (see e.g. Corollary 10.10 of [67]) that
@@
The returned matrix has slightly more than $k$ columns, often referred to as “oversampling,” but, importantly, there is no explicit SVD involved in the computation. Even in this case, floating point arithmetic already spoils the approximation guarantees: we cannot compute $\boldsymbol{\mathrm{Q}}$ without rounding errors. Stability analysis of QR factorization is typically carried out in the backward-error sense [70, 39], which makes the analysis of low-rank approximation PCA algorithms even more complicated.
@@ G.3 Bit complexity analysis of Block-Krylov PCA
We now analyze the seminal Block-Krylov iteration algorithm of [106], which is listed in Algorithm 8 for convenience (in exact arithmetic). For the floating point analysis the following methodology is used. The main result is stated in Theorem 4.2. We shall denote by $T_{\mathsf{MM}}(\boldsymbol{\mathrm{X}},q)$ the cost of multiplying $\boldsymbol{\mathrm{X}}$ with a dense matrix with $q$ columns with a numerically stable multiplication algorithm, like the one of Theorem A.2. Formally, for the rest of this section, we assume the following subroutine.
Definition G.1 ($\mathsf{MMX}$).
Let $\boldsymbol{\mathrm{X}}\in\mathbb{R}^{m\times n}$ be the input matrix of Algorithm 8. We assume a subroutine $\mathsf{MMX}(\boldsymbol{\mathrm{B}})$, which takes as input a matrix $\boldsymbol{\mathrm{B}}$ with $k$ columns. It returns a matrix $\boldsymbol{\mathrm{C}}\in\mathbb{R}^{m\times r}$ such that
$\left\|\boldsymbol{\mathrm{C}}-\boldsymbol{\mathrm{X}}\boldsymbol{\mathrm{B}}\right\|\leq\textbf{{u}}\operatorname{poly}(m,k)\|\boldsymbol{\mathrm{X}}\|\|\boldsymbol{\mathrm{B}}\|,$
in a total of $T_{\mathsf{MMX}}(k)$ floating point operations, using at most $O(\log(1/\textbf{{u}}))$ bits of precision. We also assume $\mathsf{MMX}^{\top}(\boldsymbol{\mathrm{B}})$ which approximates the product $\boldsymbol{\mathrm{X}}^{\top}\boldsymbol{\mathrm{B}}$ with the same cost and approximation bounds.
The standard inner-product based algorithm as well as a block variant of the $\mathsf{MM}$ algorithm of Theorem A.2 satisfy this definition.
Algorithm 8 Block-Krylov Iteration (Alg. 2 of [106])
@@
Given $\boldsymbol{\mathrm{G}}$ and $\boldsymbol{\mathrm{X}}$, we can construct the Krylov matrix $\boldsymbol{\mathrm{K}}$ using $O(q)$ calls to $\mathsf{MMX}(\boldsymbol{\mathrm{G}})$. The first call $\boldsymbol{\mathrm{X}}_{1}\leftarrow\mathsf{MMX}(\boldsymbol{\mathrm{G}})$ returns a matrix $\boldsymbol{\mathrm{X}}_{1}=\boldsymbol{\mathrm{X}}\boldsymbol{\mathrm{G}}+\boldsymbol{\mathrm{E}}_{1}$ where $\|\boldsymbol{\mathrm{E}}_{1}\|\leq\textbf{{u}}\cdot O(\operatorname{poly}(mk))\|\boldsymbol{\mathrm{X}}\|\|\boldsymbol{\mathrm{G}}\|\leq\textbf{{u}}\cdot O(\operatorname{poly}(mk))$, since we assumed constant norms. If we keep performing multiplications recursively to build the Krylov matrix $\|\widetilde{\boldsymbol{\mathrm{K}}}\|$, after $O(q)$ multiplications it holds that $\|\widetilde{\boldsymbol{\mathrm{K}}}-\boldsymbol{\mathrm{K}}\|\leq\textbf{{u}}\cdot O(\operatorname{poly}(mkq))$. The total cost is $O(qT_{\mathsf{MMX}}(k))$ floating point operations. If $\boldsymbol{\mathrm{X}}$ is sparse and we use the standard inner product-based algorithm, $T_{\mathsf{MMX}}(k)=O(k{\tt nnz}(\boldsymbol{\mathrm{X}}))$. If $\boldsymbol{\mathrm{X}}$ is dense and we use Theorem A.2 then $O(qmk^{\omega-2})$ floating point operations are sufficient.
@@
So far we have approximated $\boldsymbol{\mathrm{K}}$ by $\widetilde{\boldsymbol{\mathrm{K}}}=\boldsymbol{\mathrm{K}}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{K}}}$.
In the following steps we will need an approximation for the condition number of $\boldsymbol{\mathrm{K}}$. We can get such an approximation using a variant of the $\mathsf{COND}$ algorithm (Corollary E.1), which internally uses the $\mathsf{SIGMAK}$ algorithm of Proposition E.2 based on counting queries to approximate the condition number. In that case, we start with some arbitrary error $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{K}}}\|$, and we keep keep halving it until we have a sufficiently good approximation. At each iteration $t$, the algorithm executes
$O\left(qT_{\mathsf{MMX}}(k))\right)$
floating point operations to construct the Krylov matrix using $O(\operatorname{polylog}(m,q,\tfrac{1}{\epsilon_{t}})$ bits. At most $O(\log(\kappa(\boldsymbol{\mathrm{K}})))$ iterations are required. The most expensive iteration is the last one, when $\epsilon_{t}\approx 1/\kappa(\boldsymbol{\mathrm{K}})$, which gives a total cost of at most
$O\left(q\log(\kappa(\boldsymbol{\mathrm{K}}))T_{\mathsf{MMX}}(k)\right)$
floating point operations using at most $O(\operatorname{polylog}(m,q,\kappa(\boldsymbol{\mathrm{K}})))$ bits. It returns a value $\widetilde{\kappa}\in\Theta(\kappa(\boldsymbol{\mathrm{K}}))$, and it succeeds with high probability.
@@
$\textbf{{u}}\leq\epsilon^{\prime}\frac{c}{\operatorname{poly}(mq)\kappa(\boldsymbol{\mathrm{K}})},$
@@
Let $\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}=\boldsymbol{\mathrm{E}}_{1}\boldsymbol{\mathrm{\Phi}}_{2}+\boldsymbol{\mathrm{E}}_{\mathsf{QR}}$. We have that $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}\|=\|\boldsymbol{\mathrm{E}}_{1}\boldsymbol{\mathrm{\Phi}}_{2}+\boldsymbol{\mathrm{E}}_{\mathsf{QR}}\|\leq\epsilon^{\prime}+\boldsymbol{\mathrm{E}}_{\mathsf{QR}}.$ From Theorem A.6 we can ensure that $\|\boldsymbol{\mathrm{E}}_{\mathsf{QR}}\|\leq\epsilon^{\prime}$ if we use $O(\log(\tfrac{mq\kappa(\widetilde{\boldsymbol{\mathrm{K}}})}{\epsilon^{\prime}}))$ bits. Thus, $\widetilde{\boldsymbol{\mathrm{Q}}}$ is an approximate orthogonal basis for the range of the true Block-Krylov matrix $\boldsymbol{\mathrm{K}}$.
To summarize, as long as $\textbf{{u}}\leq\epsilon^{\prime}\frac{c}{\operatorname{poly}(mq)\kappa(\boldsymbol{\mathrm{K}})}$, which can be achieved by replacing $\kappa(\boldsymbol{\mathrm{K}})$ with $\widetilde{\kappa}$ above, then we can compute $\widetilde{\boldsymbol{\mathrm{Q}}}=\boldsymbol{\mathrm{Q}}\boldsymbol{\mathrm{\Phi}}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}$, where $\boldsymbol{\mathrm{Q}}$ is the true orthonormal basis from the QR factorization of $\boldsymbol{\mathrm{K}}$, $\boldsymbol{\mathrm{\Phi}}$ is a $qk\times qk$ orthogonal matrix, and $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}\|\leq\epsilon^{\prime}$. The total cost is $O(m(qk)^{\omega-1})$ floating point operations for Theorem A.6, using $O\left(\log(\tfrac{mq\kappa(\boldsymbol{\mathrm{K}})}{\epsilon^{\prime}})\right)$ bits.
@@
In Step 5, Algorithm 8 forms the matrix $\boldsymbol{\mathrm{M}}$ to compute its top-$k$ singular vectors. To analyze the computation of $\boldsymbol{\mathrm{M}}$, the first observation is that $\boldsymbol{\mathrm{Q}}$ in lines 4-7 can be replaced by any basis for the column space of $\boldsymbol{\mathrm{K}}$. In particular, we replace it by $\widetilde{\boldsymbol{\mathrm{Q}}}$. We then perform the multiplication in two steps: $\widetilde{\boldsymbol{\mathrm{M}}}_{1}^{\top}=\mathsf{MMX}^{\top}(\widetilde{\boldsymbol{\mathrm{Q}}})$, which returns $\widetilde{\boldsymbol{\mathrm{M}}}_{1}=\widetilde{\boldsymbol{\mathrm{Q}}}^{\top}\boldsymbol{\mathrm{X}}+\boldsymbol{\mathrm{E}}_{1}^{\mathsf{MM}}$, where $\|\boldsymbol{\mathrm{E}}_{1}^{\mathsf{MM}}\|\leq\textbf{{u}}\cdot O(\operatorname{poly}(qknm))\|\widetilde{\boldsymbol{\mathrm{Q}}}\|\|\boldsymbol{\mathrm{X}}\|\in\textbf{{u}}\cdot O(\operatorname{poly}(qm))$. Then we compute $\widetilde{\boldsymbol{\mathrm{M}}}\leftarrow\mathsf{MM}(\widetilde{\boldsymbol{\mathrm{M}}}_{1},\widetilde{\boldsymbol{\mathrm{M}}}_{1}^{\top})=\widetilde{\boldsymbol{\mathrm{M}}}_{1}\widetilde{\boldsymbol{\mathrm{M}}}_{1}^{\top}+\boldsymbol{\mathrm{E}}_{2}^{\mathsf{MM}}$ where $\|\boldsymbol{\mathrm{E}}_{2}^{\mathsf{MM}}\|\leq\textbf{{u}}\cdot O(\operatorname{poly}(qkn))\|\boldsymbol{\mathrm{M}}_{1}\|^{2}\in\textbf{{u}}\cdot O(\operatorname{poly}(qkn))$, where the last is implied if u is sufficiently smaller than $1/\operatorname{poly}(qm)$. Putting everything together we can write
@@
where $\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}$ contains the errors in $\widetilde{\boldsymbol{\mathrm{Q}}}$ and also the ones from the multiplication in floating point, and $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{M}}}\|\leq O(\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{Q}}}\|)\in O(\epsilon^{\prime})$. $\epsilon^{\prime}$ is as in Step 3. So far we have assumed $O(\log(\tfrac{mq\kappa(\widetilde{\boldsymbol{\mathrm{K}}})}{\epsilon^{\prime}}))$ bits.